# Fourfold path to Thermality: Inequivalent purifications of Rindler wedge

### v1

@@ Abstract
We investigate thermal behaviour in quantum fields by analysing a hierarchy of null-shifted Rindler wedges in Minkowski spacetime. Starting from the Minkowski vacuum restricted to an initial Rindler wedge, we construct several inequivalent transformation paths, including direct Minkowski-to-Rindler mappings, spatial translations, and sequential null displacements, and study the resulting particle content using Bogoliubov transformations. In the standard Unruh effect, horizon-induced entanglement between left- and right-moving sectors gives rise to Gibbsian thermality, with both sectors exhibiting thermal occupation numbers described by a mixed Gibbs state. In contrast, we show that null-shifted wedge constructions lead to a selective and non-Gibbsian form of thermality, in which only a single chiral sector develops Bose–Einstein–distributed occupation numbers. At the same time, the complementary sector remains in a vacuum. Along composite transformation paths, the global Minkowski state remains pure, and the induced states associated with null-shifted wedges are themselves pure tensor-product states. The resulting thermal behaviour arises from Bogoliubov mixing and modular time evolution rather than from entanglement across a horizon or the emergence of a Gibbs density matrix. These findings demonstrate the existence of inequivalent purifications of thermal occupation spectra and clarify the distinct roles of horizon structure, observer dependence, Bogoliubov transformations, and entanglement in the emergence of thermality. In this sense, the null-shifted construction may be viewed as a converse of the Unruh effect, wherein thermal behaviour arises without entanglement-induced mixedness. Our results establish that the appearance of thermal spectra in Rindler wedges does not uniquely imply Gibbsian thermality or horizon-induced entanglement, highlighting the operational independence of thermality and entanglement in relativistic quantum field theory.
Keywords: Unruh effect; Rindler spacetime; Bogoliubov transformations; Selective thermalization; Quantum field theory in curved spacetime; Inequivalent purifications of vacuum
@@ I Introduction
The Unruh effect reveals a deep interplay between quantum field theory, spacetime structure, and thermality. A uniformly accelerated observer in Minkowski spacetime perceives the inertial vacuum as a thermal state, with a temperature proportional to the proper acceleration. In the Rindler description, this phenomenon is understood as a consequence of horizon-induced entanglement; the Minkowski vacuum is entangled between causally disconnected Rindler wedges, and restricting the global vacuum state to a single wedge yields a mixed thermal density matrix[24]. From the perspective of algebraic quantum field theory, this thermality arises because the local algebra associated with a Rindler wedge possesses a nontrivial commutant[13], and the reduced state is obtained by tracing over the complementary algebra. An astrophysical black hole is predicted to radiate thermal Hawking radiation with a temperature roughly proportional to the inverse of the black hole mass [10]. The Rindler spacetime offers a simpler setting where black hole radiation can be studied and understood.
While the Unruh effect is traditionally formulated in this setting, it is natural to pose a converse question: given a Rindler wedge in which an accelerated observer detects a thermal spectrum, what are the possible global states and spacetime regions that can give rise to such thermality upon restriction to that wedge? The Minkowski vacuum provides a canonical realisation of this phenomenon, but it is not evident that it is unique. Addressing this converse problem allows us to investigate whether thermality in Rindler spacetime must necessarily originate from entanglement across horizons, or whether alternative mechanisms can produce thermal spectra in accelerated frames. In this work, by a thermal spectrum we mean the characteristic occupation-number distribution detected by an accelerated observer, without assuming that the underlying state is a Gibbs thermal density matrix.
In this work, we explore this question by systematically comparing two classes of transformations relating different Rindler wedges, spatial shifts and null shifts. We focus on the wedge $R_{3}$, which serves as the observer’s accessible region. In the case of spatial shifts, the transformation preserves the causal structure in a manner closely analogous to the standard Rindler decomposition. The resulting state in $R_{3}$ exhibits thermal behaviour that can still be traced to entanglement between causally separated regions, and the associated reduced state remains mixed. In this sense, spatial shifts retain the essential structural features of the Unruh effect.
The situation changes qualitatively in the presence of null shifts. A null shift corresponds to a translation along a horizon generator and, unlike the standard Rindler partition, does not introduce a new causally inaccessible region. Instead, null shifts preserve causal accessibility along the horizon generators and therefore do not lead to a decomposition of the Hilbert space into complementary algebras. As a result, the algebra associated with a null-shifted wedge does not admit a nontrivial commutant that would play the role of an inaccessible region[13, 30]. No tracing over external degrees of freedom is performed, and the global state remains pure when restricted to the null-shifted wedges $R_{2}$, $R_{4}$, and $R_{3}$. Nevertheless, the Bogoliubov transformation relating the corresponding mode decompositions gives rise to a thermal particle spectrum as detected by an accelerated observer. In particular, the states in $R_{2}$ and $R_{4}$ naturally correspond to outgoing and ingoing modes, respectively, while the state in $R_{3}$ is related to them by a null translation of the corresponding mode decomposition.
This demonstrates that the thermality observed in null-shifted wedges is not of the Unruh type. In the standard Unruh effect, thermality is inseparably tied to horizon-induced entanglement and the resulting thermofield-double structure obtained by tracing over causally inaccessible degrees of freedom. By contrast, the null-shift construction yields thermal particle spectra within globally pure states and without entanglement between distinct wedges. In this precise sense, null shifts realise a converse (in the sense of reversing the logical origin of thermality) to the Unruh effect; rather than thermality emerging from tracing over entangled degrees of freedom, it arises from kinematical Bogoliubov mixing induced by horizon-preserving translations.
We show that the converse problem admits multiple inequivalent realisations, corresponding to distinct embeddings of observer-detected thermal spectra in Rindler spacetime. By contrasting spatial and null shifts, we uncover a richer structure underlying thermality in accelerated frames and demonstrate that thermal spectra in Rindler wedges do not uniquely characterise the Unruh effect. Our results highlight that thermality in quantum field theory can arise through mechanisms fundamentally different from horizon-induced entanglement, thereby broadening the conceptual landscape of the Unruh phenomenon.
We consider a massless scalar field defined on two-dimensional Minkowski spacetime and investigate the emergence of thermality in a fixed Rindler wedge $R_{3}$. All quantum states are taken to be globally defined on Minkowski spacetime, and the physical properties observed in $R_{3}$ arise from restricting these global states to the algebra of observables associated with the wedge.
Path 1 corresponds to the standard Unruh setup, in which the global state is the Minkowski vacuum $|0_{M}\rangle$. When restricted to the Rindler wedge $R_{3}$, this state yields a reduced density matrix that is thermal with respect to the boost generator associated with $R_{3}$, reproducing the usual Unruh effect.
In Path 2, we consider a larger Rindler wedge $R_{1}$ such that $R_{3}\subset R_{1}$, with the bifurcation points of $R_{1}$ and $R_{3}$ separated by a spacelike interval. The global state is chosen so that its restriction to $R_{1}$ coincides with the Rindler vacuum associated with $R_{1}$. This state is not Hadamard when viewed as a state on Minkowski spacetime and exhibits singular behaviour on the Rindler horizon. Nevertheless, we analyse its restriction to $R_{3}$ as a useful comparison case, illustrating how thermodynamic properties in $R_{3}$ depend on the choice of global vacuum.
Paths 3 and 4 are constructed in closer analogy with black hole physics. We consider null-shifted Rindler wedges $R_{2}$ and $R_{4}$, both of which contain $R_{3}$, and global quantum states characterised by nontrivial energy fluxes along null directions. In Path 3, the global state carries an outward flux of scalar particles, analogous to outgoing radiation in the near-horizon region of a collapsing spacetime. In Path 4, the global state is characterised by an inward thermal flux of scalar particles, analogous to ingoing thermal radiation outside a black hole horizon. In both cases, the properties observed in $R_{3}$ arise from restricting these globally defined flux-carrying states to the wedge $R_{3}$.
@@
We also show an interesting effect; we consider a Rindler wedge with either an outgoing or an ingoing particle flux and convert it to a particle density in a Rindler wedge that is a subset of the previous wedge. The paths three and four in the previous paragraph use this effect to determine the particle density.
We mention a few caveats and assumptions followed in this work. The analysis is done using two methods. One is via mode functions and explicitly evaluating the Bogoliubov coefficients. The other method is using the Virasoro anomaly in two-dimensional conformal field theory (2d CFT). Since in this article we consider a massless scalar field in two dimensions, the techniques of 2d conformal field theory are applicable. We use the well known Virosoro anomaly to evaluate the stress tensor to prove the particle content in various Rindler wedges. In the analysis done in this article, we assume that the magnitude of acceleration is large and the calculations are valid close to the horizon of the uniformly accelerating observer.
As mentioned earlier, the starting point of Path 2 is Rindler spacetime with a scalar field in the Rindler vacuum. When viewed as a state on Minkowski spacetime, the Rindler vacuum is not Hadamard and exhibits a divergent local stress tensor on the Rindler horizon. As discussed in the literature [1], the Boulware vacuum in black hole spacetimes exhibits analogous near-horizon behaviour, although it is globally distinct and defined in curved spacetime. We include Path 2 as a diagnostic example rather than a physically realisable state. It is important to emphasise that the thermality encountered in null-shifted wedges is non-Gibbsian in nature, although thermal particle spectra arise, the corresponding states are not thermal density matrices of the form $\rho\propto e^{-\beta H}$, but remain globally and locally pure.
To clarify the conceptual distinctions between the different mechanisms of thermality studied in this work, we summarise their key features in Table 1. In particular, the table highlights the contrast between Gibbsian thermality arising from horizon-induced entanglement and the non-Gibbsian thermal behaviour generated by null-shifted Bogoliubov transformations.
Scenario What changes? What the observer sees in $R_{3}$ State nature Entanglement? Type of thermality
Unruh effect Accelerated observer Thermal spectrum in both sectors Mixed state Yes Gibbsian (true thermal state)
Spatial shift of wedges Spatial displacement of horizon Thermal spectrum in both sectors Mixed state Yes Gibbsian (Unruh-type)
Null shift of wedges Translation along the horizon Thermal spectrum in one sector only Pure state No Non-Gibbsian (selective)
Table 1: Physical interpretation of thermality in different Rindler constructions. The Unruh effect and spatial shifts yield genuine Gibbsian thermality associated with entanglement across horizons. In contrast, null-shifted wedges exhibit selective, non-Gibbsian thermal behaviour, where thermal particle spectra emerge without entanglement and the underlying quantum states remain pure.
The remainder of this paper is organised as follows. In Section II, we review the quantisation of a massless scalar field in two-dimensional Minkowski spacetime and introduce the relevant Rindler wedges, mode decompositions, and vacuum states that form the basis of our analysis. In Section III, we present the two complementary methods used throughout the paper to diagnose thermality: an explicit Bogoliubov transformation analysis of mode functions, and an independent computation of the stress tensor using the Virasoro anomaly in two-dimensional conformal field theory.
Section IV constitutes the core of the paper, where we investigate the particle content observed in a fixed Rindler wedge $R_{3}$ along four distinct paths. Path 1 reproduces the standard Unruh effect by restricting the Minkowski vacuum to $R_{3}$. Path 2 considers a global state that coincides with the Rindler vacuum in a larger wedge and illustrates how non-Hadamard behaviour leads to inequivalent thermodynamic properties upon restriction. Paths 3 and 4 involve null-shifted Rindler wedges and global states carrying outgoing or ingoing fluxes, respectively, and demonstrate how thermal particle spectra can arise in $R_{3}$ within globally pure states through kinematical Bogoliubov mixing.
In Section V, we compare the stress tensor expectation values obtained along different paths and show that, while particle spectra may appear thermal in all cases, the underlying states are physically inequivalent. Finally, in Section VI, we discuss the broader implications of our results for the interpretation of horizon thermality, clarify the sense in which null-shift constructions realise a converse to the Unruh effect, and comment on possible connections to black hole physics.
@@ II The Problem Statement and SetUp
We consider a fixed Rindler wedge $R_{3}$ and a massless real scalar field defined globally on two-dimensional Minkowski spacetime. We aim to investigate how thermal particle spectra observed in $R_{3}$ can arise from restricting globally defined quantum states to the algebra of observables associated with $R_{3}$. In particular, we ask whether the global spacetime region and quantum state leading to a given particle content in $R_{3}$ are uniquely determined. As outlined in the Introduction, we show that this is not the case; distinct choices of global states and spacetime regions, related by spatial or null shifts of Rindler wedges, can lead to identical or similar thermal spectra when restricted to $R_{3}$. To address this question systematically, we introduce below several distinct paths corresponding to different sequences of wedge restrictions and coordinate transformations.
@@ II.1 Four inequivalent paths
We discuss below the various paths leading to a common end point viz wedge $R_{3}$ having a thermal distribution of particles in the high acceleration regime (close to the horizons). As stated before in the introduction, we show that various starting points yield the common endpoint. By starting points, we mean various spacetimes with the massless scalar fields in corresponding particle states. We demonstrate this by explicitly choosing the starting points and evaluating the final particle content in $R_{3}$ along the given path.
II.1.1 Path1: Minkowski (in vacuum state) MMto Rindler wedge $\mathrm{R_{3}}$
In this path, we consider Minkowski spacetime with the massless scalar field in a vacuum state. The wedge $R_{3}$ is located such that its bifurcation point is situated to the right of the origin. The Unruh effect states that the reduced state of the massless scalar field in $R_{3}$ is a thermal distribution of particles.
@@ II.1.2 Path 2: Rindler wedge $R_{1}$ in vacuum state to Rindler wedge $R_{3}$ via Spatial Shift
In Path 2, the wedge $R_{3}$ is defined from $R_{1}$ by a spatial translation, characterised by displacement parameter $\Delta_{1}$ (cf. Figure 1). Although the bifurcation point of $R_{3}$ is shown along the $x_{1}$ axis, its precise location may be generalised to any point within the $(u,v)$ light-cone plane.
II.1.3 Path 3: Sequential Null Shifts
Path 3 traces a sequence $R_{1}\rightarrow R_{2}\rightarrow R_{3}$, where
$R_{2}$ is obtained from $R_{1}$ by displacing its origin along the future-directed null $V_{1}$-axis by a Minkowski parameter $\Delta_{2}$, $R_{3}$ is then constructed by a further displacement of $R_{2}$ along the null $U_{2}$-axis by parameter $\Delta_{3}$ (see Figure 1). For illustration, Figure 1 depicts the case where $\Delta_{2}=\Delta_{3}$, but this equality is not a physical constraint and can be ensured by an appropriate Lorentz transformation, since these parameters are not Lorentz scalars.

### v2

# Four inequivalent paths to Thermality in Minkowski spacetime

@@ Abstract
In this article we explore the answer to an inverse question to Unruh effect. Given a two dimensional Rindler wedge $R$ with a thermal distribution of massless scalar particles, we explore the supersets of $R$, whose reduced state yields the observed particle content in $R$. If we restrict our analysis to particle content close to the Rindler horizon in $R$, we show that the answer to the inverse question is not unique. The analysis is done using two methods, first uses Bogolibov coefficients and the second method uses Virasoro anomaly. We identify four inequivalent paths with various ”parent” spacetimes of $R$ from which we can arrive at the given particle content in $R$. We show that the supersets are: Minkowski spacetime, a Rindler wedge in vacuum, a Rindler wedge with a thermal flux of left moving particles and lastly a Rindler wedge with a thermal flux of right moving flux of particles. We also show an interesting phenomenon: flux to density conversion, wherein a thermal flux of either left moving or right moving particles can be converted to thermal density of particles involving both left moving and right moving particles. Based on the analysis done in the article, we present qualitative arguments regarding the possibility that the continuous evaporation of a black hole might be punctuated by a series of pauses and bursts of radiation.
@@ I Introduction
The Unruh Davies effect is a well-studied and well-understood phenomenon. When we consider a two-dimensional Minkowski spacetime with a quantum field in its vacuum state, a uniformly accelerating observer perceives the field in vacuum as an excited state with a thermal distribution of particles [20]. A Rindler spacetime is a static spacetime constructed by considering a continuum of accelerating observers. It is a subset of Minkowski spacetime, and when a Minkowski spacetime contains a quantum field in a vacuum state, it appears as a thermal state in Rindler spacetime. A related idea is that of Hawking radiation. An astrophysical black hole is predicted to radiate a thermal flux of radiation known as Hawking radiation with a temperature roughly proportional to the inverse of black hole mass [8]. Rindler spacetime offers a simpler setting in which black hole radiation can be studied and understood.
Nested and shifted Rindler wedges provide a rigorous geometric framework for exploring observer-dependent thermodynamics and subsystem entanglement in quantum field theory. Exploring the kinematic consequences of nesting, Lochan and Padmanabhan showed that sequences of spatially shifted wedges exhibit “relic thermality”: the vacuum of an enveloping wedge appears universally thermal to any strictly nested observer, regardless of the spatial shift’s magnitude [10]. A shifted Rindler model was also used in Ref. [7], who demonstrated that tracing out degrees of freedom across nested regions (e.g., $R_{2}\subset R_{1}$) can be path-dependent. Furthermore, extensions to null-shifted wedges by the authors in Ref. [12] revealed “selective particle excitations,” a geometric consequence where only specific momentum modes or chiralities are thermally excited. Ultimately, these works demonstrate that the precise causal nesting of Rindler wedges can be used as toy models to study dynamical evolution of black hole apparent horizon.
In this article, we are interested in the converse question to the Unruh effect. We have a Rindler wedge $R_{3}$ with a massless scalar field with particle distribution in a thermal state. What are the supersets $\mathcal{A}$ of $R_{3}$ ($R\subset A$) whose reduced state yields a thermal state in $R_{3}$? Stated differently, if we have a uniformly accelerating observer in Minkowski spacetime who detects a thermal bath of scalar particles, what can he/she deduce about the spacetime beyond his/her horizon? Since he/she is a uniformly accelerating observer, he/she has a future and past horizon, and he/she is therefore restricted to a Rindler wedge. We ask a converse question. What are the possible supersets of his/her Rindler wedge whose reduced state yields the observed particle distribution? One obvious answer is, of course, Minkowski spacetime with the scalar field in vacuum. But is the answer unique? We justify in this paper that the answer is not unique by showing that thermality in a given Rindler wedge can occur along four inequivalent paths; therefore, we argue that the converse question has multiple answers.
Path one is as stated before, is from a Minkowski spacetime with the scalar field in a vacuum state. Path two involves a Rindler wedge $R_{1}$ with the scalar field in vacuum (the Rindler wedge $R_{1}$ is located such that $R_{3}\subset R_{1}$, such that the bifurcation point of $R_{1}$ and $R_{3}$ are separated by a spacelike interval. Path three starts with a Rindler wedge $R_{2}$ (with $R_{3}\subset R_{2}$) with an inward flux of scalar particles. $R_{3}$ is a wedge shifted from $R_{2}$ along a null ray as explained in later sections. Path four starts with a Rindler wedge $R_{4}$ again (with $R_{3}\subset R_{4}$), and in contrast with path three, there is an outward thermal flux of scalar particles. $R_{3}$ is a wedge shifted along a null ray from $R_{4}$ as illustrated in later sections.
@@
We also show a new interesting effect: flux-to-density conversion. We consider a Rindler wedge with either an outgoing or an ingoing particle flux and show how to convert it to a particle density in a Rindler wedge that is a subset of the previous wedge. The paths three and four in the previous paragraph illustrate this effect.
We mention a few aspects and assumptions followed in this work. The analysis is done using two methods. One is via mode functions and the explicit evaluation of the Bogoliubov coefficients. The Bogoliubov transformation approach operates at the level of mode expansions, mixing creation and annihilation operators to relate distinct vacua and thereby connect asymptotic particle states. The other method uses the Virasoro anomaly in two-dimensional conformal field theory (2D CFT). Since we consider a massless scalar field in 2D for this work, the techniques of 2D conformal field theory are applicable. We use the Virasoro anomaly to evaluate the stress tensor and thereby determine the particle content in various Rindler wedges. By contrast to the Bogoliubov approach, the stress tensor and conformal anomaly method evaluate renormalized local operator expectation values, such as energy fluxes, and reveal how anomalies encode radiation and particle production. While the Bogoliubov method emphasizes the relation between asymptotic states, the stress tensor method provides a local field‑theoretic account of fluxes and anomalies. Both approaches ultimately describe the same physical phenomenon, but from different calculational perspectives. In the analysis presented in this article, we take the assumption that the frequencies involved are large and that the calculations are applicable near the horizon of the uniformly accelerating observer.
As mentioned before, the starting point of Path-2 is the Rindler spacetime with a scalar field in the Rindler vacuum. Now, the Rindler vacuum is not a physically well-defined state in the context of Minkowski spacetime, since the local stress tensor with respect to the Rindler vacuum diverges on the horizon. As demonstrated in past literature [1], the Boulware vacuum (in the near-horizon limit, the Boulware vacuum behaves like the Rindler vacuum, as the Boulware vacuum is defined in curved spacetime and globally differs from the Rindler vacuum, whereas the Rindler vacuum lives in flat spacetime and only covers part of it) in the context of black holes. The Rindler spacetime and the analysis done in the article serve as a toy model for the realistic case involving evolving black holes and the possible signatures of the evolving apparent horizons. As pointed out earlier, the Rindler spacetime maps to the Boulware vacuum in curved spacetime. As we discuss in the conclusions section of the article, the Rindler vacuum in the toy model and the corresponding Boulware vacuum in the curved spacetime setting may play a role in the context of black hole evaporation.
@@ II The Problem Statement and SetUp
Suppose we have a Rindler wedge $R_{3}$ in which we consider a real massless scalar field with the particle content in a thermal distribution. We are interested in the supersets $A$ of $R_{3}$ that yield the observed particle content to be thermal upon finding the reduced state in $R_{3}$ from $A$. As stated in the introduction, we show that $A$ is not unique. We begin by formulating the setup necessary for the analysis and subsequently present the different possible paths.
@@ II.1 Four inequivalent paths
We discuss below the various paths leading to a common endpoint, viz., the wedge $R_{3}$ with a thermal distribution of particles in the high-acceleration regime (close to the horizons). As stated before in the introduction, we show that various starting points yield the common endpoint. By ‘starting points,’ we mean various spacetimes with massless scalar fields in their corresponding particle states. We demonstrate this by explicitly choosing the starting points and evaluating the final particle content in $R_{3}$ along the given paths. The coordinate definitions are given below, Two dimensional Minkowski spacetime is denoted as $\mathcal{M}$, parametrized by inertial coordinates $(X,T)$, and a hierarchy of Rindler wedges [16], denoted $(R_{i})$, each described by coordinates $(x_{i},t_{i})$ for $i\in\{1,2,3,4\}$. As depicted in Figs . 1 and  2, the nested structure of Rindler wedges are understood as inclusions of spacetime regions determined by their null horizons, such that $R_{3}\subset R_{2}\subset R_{1}\subset M$, and $R_{3}\subset R_{4}\subset R_{1}\subset M$, with individual wedges distinguished by the shaded regions in pale red, blue, and green. Wedges $R_{2}$ and $R_{3}$ are generated from $R_{1}$ by successive null displacements along the relevant Minkowski light rays, a construction central to our analysis.
II.1.1 Path 1: Minkowski (in vacuum state) $M$ to Rindler wedge $\mathrm{R_{3}}$
In this path, we consider Minkowski spacetime with the real massless scalar field in a vacuum state. The wedge $R_{3}$ is located such that its bifurcation point is situated to the right of the origin. The Unruh effect states that the reduced state of the massless scalar field in $R_{3}$ is a thermal distribution of particles. Keeping in perspective the method followed in the article, we note that we can also arrive at this result by subdividing this path into two sub-steps. We note that the metric in coordinates $T,X$ is $ds^{2}=-dT^{2}+dX^{2}$. We choose null rays as U=TXU=T-X and $V=T+X$. The two sub-paths are stated below: a) We divide the Minkowski spacetime into two parts by choosing coordinate patches $U,V$ with a new map $u,V$ with $u$ given by $U=-e^{-au}/a-\delta$ where $a$ is an arbitrary real positive parameter. We note that the null coordinates range over $\infty<U<\infty$ and, similarly, $\infty<u<\infty$. We note that when coordinate $u$ goes from -$\infty$ to $\infty$, the coordinate $U$ ranges from 0 to $\infty$. This is defined as half-sided modular inclusion [26]. b) We now go from half of Minkowski spacetime to one fourth by defining one more coordinate transformation given by $u,v$ where $u$ is the same as in subpart a and $V=e^{av}/a+\gamma$ with $\gamma$ a positive constant. We can see that this gives us the Rindler right wedge, provided we note the restriction that $U>0$ and $V<0$. We note that we could interchange the order of subparts $a$ and $b$. In the standard Unruh effect, it is well known that this ordering does not make a difference to the final particle content in the Rindler wedge.
@@ II.1.2 Path 2: Rindler wedge $R_{1}$ in vacuum state to Rindler wedge $R_{3}$ via Spatial Shift
In Path 2, the wedge $R_{3}$ is defined from $R_{1}$ by a spatial translation, characterized by displacement parameter $\Delta_{1}$ (cf. Figure 1). Although the bifurcation point of $R_{3}$ is shown along the $x_{1}$ axis, its precise location may be generalized to any point within the $R_{1}$ wedge. Though the answer to this path is worked out in Refs. [10, 7] using the method of Bogoliubov coefficients, we present an alternative analysis here using the Virasoro anomaly approach. We note that in a given Rindler wedge, say for example in $R_{1}$, the magnitude of the acceleration of an observer progressively decreases as one moves to the right along the $t_{1}=0$ line ($x_{1}-axis$). This acceleration tends to zero as one approaches $x_{1}->\infty$. The same is true for an observer in $R_{3}$ as one moves to the right ($x_{3}->\infty$). It is reasonable to expect that any particle excitations in the wedge $R_{3}$ are more prominent close to the horizons. We demonstrate this in the article.
II.1.3 Path 3: Null Shifts
Consider a Rindler wedge $R_{2}$ with the following particle content. The left-moving particles are present with a thermal distribution, while there are no right-moving particles. There is therefore a net left-moving thermal flux of particles. We show that starting from $R_{2}$ with the above-mentioned particle distribution, we arrive at a thermal distribution of particles in $R_{3}$ with both the left-moving and right-moving particles with a thermal distribution of particles with the standard Rindler temperature. We note that this is true close to the horizon.
An alternative definition of path 3 traces a sequence $R_{1}\rightarrow R_{2}\rightarrow R_{3}$, where $R_{2}$ is obtained from $R_{1}$ by displacing its origin along the future-directed null $V_{1}$-axis by a Minkowski parameter $\Delta_{2}$, $R_{3}$ is then constructed by a further displacement of $R_{2}$ along the null $U_{2}$-axis by parameter $\Delta_{3}$ (see Figure 1). For illustration, Figure 1 depicts the case where $\Delta_{2}=\Delta_{3}$, but this equality is not a physical constraint and can be ensured by an appropriate Lorentz transformation, since these parameters are not Lorentz scalars. In our article, we present the latter definition of path 3, where the starting point is again a vacuum in wedge $R_{1}$.